OpenStudy (anonymous):

Penelope is making a quilt that has 4 sections, each of which is 2 feet in length. What is the length of the entire quilt?

OpenStudy (anonymous):

P=2X8 = 16 Since quilt are always sold in rectangular form

OpenStudy (anonymous):

Thank u :D

OpenStudy (anonymous):

Penelope is making a quilt that has 4 sections, each of which is 2 5/8 feet in length. What is the length of the entire quilt

OpenStudy (anonymous):

2 5/8 = 21/8 P= 21/8 * 8 = 168/8 = 21 Whole

OpenStudy (anonymous):

my choices are A.2 3/8 B.6 5/8 C. 7 7/8 D. 10 1/2

OpenStudy (anonymous):

10 1/2 P= 21/8 * 4 = 10 1/2 = 10 Whole 1/2

OpenStudy (anonymous):

Your question is very vague the questtion setter may have set wrongly

OpenStudy (anonymous):

Thank u